# Nearly Zero-Cost Protection Against Mimicry by Personalized Diffusion Models

### v1

@@ 4 Experiment
Implementation Details. We set the number of perturbations $K$ in MoP as four. Other hyperparameters and details on training and inference are described in Suppl.
Datasets. We utilize four domains when constructing both the training and benchmark datasets: object, face, painting, and cartoon. These multiple scenarios allow us to analyze the models’ protection performance more closely to real applications, as images from various domains can be encountered. The dataset details are provided in Suppl.
Baselines. We compare with existing diffusion-based image protection frameworks: AdvDM [liang2023adversarial], PhotoGuard [salman2023raising], Anti-DB [van2023anti], Mist [liang2023mist], Impasto [ahn2024imperceptible], and Diff-Protect (SDST) [xue2023toward]. Except ours, all other methods rely on iterative optimization, using texture or semantic losses.
Evaluation. Latency is measured on a M1 Max CPU and an A100 GPU. To evaluate invisibility, we use DISTS [ding2020image], and FID [heusel2017gans] is utilized to measure the protection efficacy. In Suppl. we provide additional metrics for a more comprehensive comparison. Note that previous studies typically fix the protection strength across models for comparison, but we found that protection efficacy and invisibility vary significantly depending on the protection loss used. For instance, PhotoGuard [salman2023raising] and AdvDM [liang2023adversarial] exhibit different trends since they use disparate objectives. In addition, evaluating protection efficacy only without invisibility leads to a narrow assessment that might overlook real-world needs. To address these, we adjust the protection strength to match the protection level across methods to ensure a fairer comparison and better reflection of practical requirements. However, one might also be curious about how the performance appears when the protection strength is equalized across all methods. Hence, we also provide evaluation with a fixed strength in Suppl.
@@ 4.1 Model Comparison
As shown in Figure 1(a) and Table 2, FastProtect is ultra-fast; 125$\times$ faster on CPU and 175$\times$ faster on GPU compared to the second fastest model, PhotoGuard [salman2023raising]. Remarkably, FastProtect maintains consistent latency even as the input image size increases, achieving near real-time performance even for $2048^{2}$-px images. In contrast, all other frameworks show exponential increases in latency (Figure 1(a)), which poses a significant issue in the real world where many recent artworks are high-resolution. The unprecedented speed of our model makes it highly user-friendly.
@@
$\rightarrow$ SD-v2.1 [rombach2022high] $\rightarrow$ SD-XL [podell2023sdxl] TI [gal2022image] DreamStyler [ahn2023dreamstyler]
@@
FastProtect also demonstrates a superior trade-off between protection efficacy and invisibility compared to other frameworks (Figure 1(b)). In comparisons where the protection strength is adjusted to yield similar protection efficacy (Table 2), FastProtect consistently achieves the best invisibility across most protection domains, with the second-best performance in the painting domain. In our analysis, FastProtect shows the strong advantage in invisibility within areas that have flat textures. These features result in significant improvements in the object and cartoon domains. For example, Figure 5 shows that in the first row, the sky region is rendered nearly invisible by FastProtect, and the same applies to the cartoon in the second row. As demonstrated in Suppl. even when we fix the protection strength of all methods to the same value, our method achieves a favorable balance in this trade-off.
@@ 4.2 Model Analysis
Ablation Study. In Table 4, we depict the component analysis of FastProtect. It is notable that UAP [moosavi2017universal] significantly deteriorates protection efficacy compared to the iterative optimization baseline, PhotoGuard [salman2023raising]. On the other hand, when we introduce the proposed MoP, it recovers the lost performance. However, if the assignment function $\mathcal{A}$ is not used, the improvement is limited, highlighting the importance of achieving a semi-image-specific nature. As we discussed, the limitations of UAP arise not only from insufficient capacity but also from its image-agnostic nature. Simply increasing the number of perturbations (i.e. MoP w/o $\mathcal{A}$) resolves the first limitation and improves performance over UAP. However, this still does not overcome the second one, which causes it to underperform compared to PhotoGuard. When we introduce the assignment function into MoP, it can dynamically account for image features. This allows MoP to match the performance of iterative optimization methods while remaining significantly faster.
@@
Robustness. Here, we analyze the vulnerability of protected images to countermeasures (Table 4). We use PhotoGuard [salman2023raising] as our baseline due to its similarity to our approach. To assess robustness, we evaluate against countermeasures such as Gaussian noise and JPEG compression. Additionally, we account for a more realistic scenario where the input image can have arbitrary dimensions. Most previous evaluations in image protection tasks assume a fixed image resolution, typically $512^{2}$, which represents an idealized scenario. In practice, image sizes and aspect ratios vary. Our model demonstrates comparable performance against these countermeasures, and notably, in the arbitrary size scenario, FastProtect performs even better. This suggests that simply using interpolation to adapt MoP to different sizes is both straightforward and effective.
Black-Box Scenario. Table 5 presents the model analysis in black-box scenarios. We examine two cases: the unknown model scenario, where we transfer protection methods to other Stable Diffusion (SD) backbones, specifically SD v2.1 [rombach2022high] and SD-XL [podell2023sdxl]. We also assess image protection against unknown personalization by applying Textual Inversion (TI) [gal2022image] and DreamStyler [ahn2023dreamstyler] to replicate the image. In these scenarios, FastProtect demonstrates superior invisibility compared to the baseline while achieving comparable or better performance in black-box settings.
@@
FastProtect + Iterative Optimization. FastProtect is extremely fast, so users with sufficient computing resources might want to invest additional computation to achieve even more effective protection. Considering this scenario, in Figure 7(c), we use the results of FastProtect as the initial perturbation and further refine them using PGD (e.g. PhotoGuard [salman2023raising]). Note that we display the baseline, PhotoGuard, with a slightly larger budget since applying PGD to FastProtect tends to decrease invisibility. Hence, the baseline budget is adjusted to match the invisibility level of FastProtect + PGD. Surprisingly, our method serves as a superior initial checkpoint for the iterative optimization techniques. For example, the baseline, requires 100 steps to match our initial results, whereas FastProtect + PGD converges in just 25 steps with a significantly higher protection efficacy.
@@ Appendix A Implementation Details
Training. We set the number of perturbations $K$ in MoP to four. When calculating MLP loss in Eq. 5, we use intermediate features from [“down_1”, “down_2”, “down_3”, “mid_0”] layers in SD VAE, and set $\lambda$ to $3.5\times 10^{-5}$. We train FastProtect on a single A100 80GB GPU in an end-to-end manner except the assignment function $\mathcal{A}$. We use a batch size of 16, using the Adam optimizer [kingma2014adam] for 40k steps with learning rate of 0.0002, and betas of (0.5, 0.99). When training, the required VRAM and compute time are around 50GB and 12 hours, respectively. Note that since we update the perturbations with Adam instead of the direct optimization through PGD, we remove the minus sign in the training loss (Eq. 5) as: $\mathcal{L}_{T}(\mathbf{x})=||\mathbf{z}-\mathbf{z}_{y}||^{2}_{2}+\frac{\lambda}{L}\sum_{l=1}^{L}||\mathcal{F}^{l}-\mathcal{F}^{l}_{y}||^{2}_{2}.$ We utilize three patterned target images, each representing low, mid, and high pattern repetition, as illustrated in Figure 8. To implement adaptive targeted protection, for each protection strength (budget) $\eta$, we obtain three MoP models (low, mid, and high), each corresponding to one of the target images. The training algorithm is detailed in Algorithm 1; note that for simplicity, it is demonstrated with a batch size of one, but increasing the batch size is straightforward.
@@ Appendix B Related Work
Universal Adversarial Perturbation (UAP). moosavi2017universal demonstrated that a single perturbation could be used to attack various images, in contrast to the traditional image-specific perturbations. This approach is significantly more efficient in terms of computation time compared to per-instance adversarial attacks because it does not require iterative optimization for each individual image. Subsequently, hayes2018learning presented a method for generating UAPs using generative models, showing improved efficiency and effectiveness over previous methods. Similarly, mopuri2018nag utilized generative networks to create adversarial examples. In addition, there have been various attempts to find UAPs tailored to the image recognition tasks [chaubey2020universal, liu2019universal, mopuri2018generalizable, mopuri2018ask]. However, these methods are not aligned with and cannot be adapted to the image protection task. Instead, in our study, we designed MoP with a focus on creating a protection framework that is both fast and maintains high performance. Our approach is specialized for image protection, differing from the general UAP-related methodologies introduced in adversarial attack field.
@@
AdvDM [liang2023adversarial] 1210s /   35s 0.129 / 199.5 0.152 / 303.7 0.117 / 346.8 0.194 / 196.5
PhotoGuard [salman2023raising]   370s /    7s 0.184 / 211.9 0.258 / 371.1 0.165 / 377.4 0.221 / 227.6
Anti-DB [van2023anti] 7278s / 225s 0.164 / 197.4 0.193 / 317.4 0.143 / 343.9 0.231 / 184.6
Mist [liang2023mist] 1440s /   40s 0.185 / 217.2 0.259 / 365.8 0.166 / 386.2 0.223 / 223.7
Impasto [ahn2024imperceptible]   830s /   19s 0.131 / 188.9 0.198 / 298.4 0.123 / 352.4 0.179 / 201.1
SDST [xue2023toward] 1410s /   24s 0.170 / 198.0 0.244 / 302.1 0.152 / 354.1 0.199 / 196.7
@@
Invisible Adversarial Perturbation. Achieving high invisibility in adversarial attacks has garnered significant research interest. Several methods focus on restricting the regions of perturbations. Some approaches use the $L_{0}$ norm to generate sparse perturbations [modas2019sparsefool, croce2019sparse], while others confine perturbations to small, salient areas to reduce visual distortions [dai2023saliency]. Various studies have targeted specific elements such as low-frequency components [guo2018low, luo2022frequency] and have utilized advanced constraints like color components [zhao2020towards, shamsabadi2020colorfool] or quality assessments [wang2021demiguise]. Additionally, luo2018towards explored techniques for creating invisible and robust adversarial examples based on the human visual system. More recently, laidlaw2020perceptual investigated perceptual adversarial robustness and adopted LPIPS [zhang2018unreasonable] in adversarial perturbation optimization.
Although invisibility has been highly emphasized in adversarial attacks, it is less explored in the image protection task. Many methods enhance invisibility implicitly by minimizing the budget used, but this often leads to a trade-off with protection efficacy, resulting in decreased performance. ahn2024imperceptible constrain perturbations based on the human visual system using various modules during the optimization process. While effective, they require many ad-hoc components. In contrast, FastProtect uses LPIPS to create perceptual maps and apply masking, which is a simpler and more streamlined approach. This differs from zhang2018unreasonable, who use LPIPS in the optimization process.
@@ Appendix C Experimental Setups
Datasets. We utilize the Object, Face, Painting, and Cartoon domains in both the training and benchmark datasets. When constructing the training dataset, we randomly sample 20k images from ImageNet [deng2009imagenet] for the Object domain, resizing all images to $512^{2}$. No augmentation is applied. For the Face domain, we use 20k images randomly sampled from the FFHQ dataset [karras2019style]. The images, originally $1024^{2}$ and face-aligned [kazemi2014one], are all downscaled to $512^{2}$, without further augmentations. The Painting dataset is created by randomly sampling 20k images from the WikiArt dataset [saleh2015large], resizing them to $512^{2}$. We filter out images with resolutions outside the range of 512 to 2048-px. For the Cartoon dataset, we use Webtoon artworks published on NAVER Webtoon, sampling 20k images only the works with permission from the original creators for research purposes. Similar to the Painting dataset, we apply resolution filtering and resize the images to $512^{2}$.
To make the benchmark dataset, we select 20 objects from the personalization subject dataset proposed by ruiz2023dreambooth for the Object domain. Each object consists of 5-6 images. When we conduct the main comparison (Table 2), we use images resized to $512^{2}$, and for the arbitrary image analysis (Table 4), we apply protection to the original resolution images. For the Face domain, we followe van2023anti and randomly sample 20 identities from VGGFace2 [cao2018vggface2], each identity consisting of 12 images resized to $512^{2}$. The Painting dataset is compiled by randomly sampling (but not overlap to the training dataset) 20 artists from the WikiArt dataset, each artist represented by 10 artworks, all resized to $512^{2}$. For the Cartoon domain, we randomly sample 20 Webtoon works (but also no overlap to the training dataset) from the Webtoon dataset. In this dataset, we only use facial images of major cartoon characters by cropping and aligning all the images. This reflects the common practice in cartoons and illustrations where characters are the main focus to both readers and artists. In the main comparison we use $512^{2}$ images, while the arbitrary analysis maintains the original resolution.
@@
AdvDM [liang2023adversarial] 0.197 0.362 0.540 220.0 0.458 2.139
PhotoGuard [salman2023raising] 0.203 0.347 0.575 223.0 0.506 2.396
Mist [liang2023mist] 0.185 0.322 0.578 217.2 0.523 2.328
SDST [xue2023toward] 0.242 0.402 0.575 219.2 0.542 2.442
Anti-DB [van2023anti] 0.239 0.413 0.510 214.4 0.439 2.148
Impasto [ahn2024imperceptible] 0.201 0.378 0.588 213.8 0.473 2.183
@@
AdvDM [liang2023adversarial] 0.173 0.338 0.528 303.8 0.466 2.493
PhotoGuard [salman2023raising] 0.189 0.357 0.557 308.7 0.630 3.113
Mist [liang2023mist] 0.154 0.310 0.560 307.5 0.667 3.131
SDST [xue2023toward] 0.244 0.431 0.561 302.1 0.629 3.019
Anti-DB [van2023anti] 0.162 0.309 0.543 301.4 0.484 2.411
Impasto [ahn2024imperceptible] 0.198 0.388 0.562 298.4 0.565 2.939
@@
Baselines. FastProtect is compared with existing diffusion-based mimicry protection frameworks: AdvDM [liang2023adversarial], PhotoGuard [salman2023raising], Anti-DreamBooth (Anti-DB) [van2023anti], Mist [liang2023mist], Impasto [ahn2024imperceptible], and SDST [xue2023toward]. We follow the official settings for all the baselines by using their official codes. However, for PhotoGuard, we use the target image of Mist [liang2023mist] by following the protocols of xue2023toward while we use the Impasto’s target image for the Impasto.
Evaluation. For the primary evaluation assessment, we use DISTS [ding2020image] to measure the invisibility of the protected image and FID [heusel2017gans] to evaluate the protection efficacy of the personalized diffusion methods. In addition, in here, we also compare the protection frameworks using other metrics: To evaluate invisibility, we include LPIPSVGG [zhang2018unreasonable] and AHIQ [lao2022attentions]. Note that LPIPSVGG measures the perceptual distance in the VGG feature space, which is different from the feature space used in our perceptual map creation module (i.e. AlexNet). For evaluating protection efficacy, we use TOPIQ-NR [chen2024topiq] and QAlign [wu2023q].
When preparing mimicry outputs, we personalize the diffusion models using the protected image with LoRA [hu2021lora] by default, employing default settings for all the personalization methods. To generate outputs, we use different inference prompts than those used during training to simulate black-box caption scenarios [van2023anti]. For example, we train the diffusion models with LoRA using captions of “A painting in <sks>  style” prompt and perform inference with “A painting of a house in <sks>  style”.

### v2

@@ 4 Experiment
Implementation Details. We set the number of perturbations $K$ in MoP as four. Other hyperparameters and details on training and inference are described in Appendix.
Datasets. We utilize four domains when constructing both the training and benchmark datasets: object, face, painting, and cartoon. These multiple scenarios allow us to analyze the models’ protection performance more closely to real applications, as images from various domains can be encountered. The dataset details are provided in Appendix.
Baselines. We compare with existing diffusion-based image protection frameworks: AdvDM [28], PhotoGuard [46], Anti-DB [51], Mist [27], Impasto [1], and Diff-Protect (SDST) [56]. Except ours, all other methods rely on iterative optimization, using texture or semantic losses.
Evaluation. The protection efficacy is evaluated against LoRA [17] on Stable Diffusion v1.5. Latency is measured using an M1 Max CPU and an A100 GPU. We employ DISTS [10] to evaluate invisibility, and FID [15] to quantify protection efficacy. In Appendix, we provide additional metrics for a more comprehensive comparison. Note that previous studies typically fix the protection strength across models for comparison, but we found that protection efficacy and invisibility vary significantly depending on the protection loss used. For instance, PhotoGuard [46] and AdvDM [28] exhibit different trends since they use disparate objectives. Moreover, evaluating protection efficacy only without invisibility leads to a narrow assessment that might overlook real-world needs. To address these, we adjust the protection strength to match the protection level across methods to ensure a fairer comparison and better reflection of practical requirements. However, one might also be curious about how the performance appears when the protection strength is equalized across all methods. Hence, we provide evaluation with a fixed strength in Appendix.
@@ 4.1 Model Comparison
As shown in Figure 1(a) and Table 2, FastProtect is ultra-fast; 125$\times$ faster on CPU and 175$\times$ faster on GPU compared to the second fastest model, PhotoGuard [46]. Remarkably, FastProtect maintains consistent latency even as the input image size increases, achieving near real-time performance even for $2048^{2}$-px images. In contrast, all other frameworks show exponential increases in latency (Figure 1(a)), which poses a significant issue in the real world where many recent artworks are high-resolution. In addition, FastProtect occupies only 1.7GB of VRAM during inference, unlike other methods that require more than 8GB. The unprecedented speed of our model makes it highly user-friendly.
@@
$\rightarrow$ SD-v2.1 [43] $\rightarrow$ SD-XL [41] TI [11] DreamStyler [2]
@@
FastProtect also demonstrates a superior trade-off between protection efficacy and invisibility (Figure 1(b)). In comparisons where the protection strength is adjusted to yield similar protection efficacy (Table 2), FastProtect consistently achieves the best invisibility across most protection domains, with the second-best performance in the painting domain. In our analysis, FastProtect shows the strong advantage in invisibility within areas that have flat textures. These features result in significant improvements in the object and cartoon domains. For example, Figure 5 shows that in the first row, the sky region is rendered nearly invisible by FastProtect, and the same applies to the cartoon in the second row. As demonstrated in Appendix, even when we fix the protection strength of all methods to the same value, our method achieves a favorable balance in this trade-off.
@@ 4.2 Model Analysis
Ablation Study. In Table 4, we depict the component analysis of FastProtect. It is notable that UAP [36] significantly deteriorates protection efficacy compared to the iterative optimization baseline, PhotoGuard [46]. On the other hand, when we introduce the proposed MoP, it recovers the lost performance. However, if the assignment function $\mathcal{A}$ is not used, the improvement is limited, highlighting the importance of achieving a semi-image-specific nature. As we discussed, the limitations of UAP arise not only from insufficient capacity but also from its image-agnostic nature. Simply increasing the number of perturbations (i.e. MoP w/o $\mathcal{A}$) resolves the first limitation and improves performance over UAP. However, this still does not overcome the second one, which causes it to underperform compared to PhotoGuard. When we introduce the assignment function into MoP, it can dynamically account for image features. This allows MoP to match the performance of iterative optimization methods while remaining significantly faster.
@@
Robustness. Here, we analyze the vulnerability of protected images to countermeasures (Table 4). We use PhotoGuard [46] as our baseline due to its similarity to our approach. To assess robustness, we evaluate against countermeasures such as Gaussian noise and JPEG compression. Additionally, we account for a more realistic scenario where the input image can have arbitrary dimensions. Most previous evaluations in image protection tasks assume a fixed image resolution, typically $512^{2}$, which represents an idealized scenario. In practice, image sizes and aspect ratios vary. Our model demonstrates comparable performance against these countermeasures, and notably, in the arbitrary size scenario, FastProtect performs even better. This suggests that simply using interpolation to adapt MoP to different sizes is both straightforward and effective.
Black-Box Scenario. Table 5 presents the model analysis in black-box scenarios. We examine two cases: the unknown model scenario, where we transfer protection methods to other Stable Diffusion (SD) backbones, specifically SD v2.1 [43] and SD-XL [41]. We also assess image protection against unknown personalization by applying Textual Inversion (TI) [11] and DreamStyler [2] to replicate the image. In these scenarios, FastProtect demonstrates superior invisibility compared to the baseline while achieving comparable or better performance in black-box settings.
@@
FastProtect + Iterative Optimization. FastProtect is extremely fast, so users with sufficient computing resources might want to invest additional computation to achieve even more effective protection. Considering this scenario, in Figure 7(c), we use the results of FastProtect as the initial perturbation and further refine them using PGD (e.g. PhotoGuard [46]). Note that we display the baseline, PhotoGuard, with a slightly larger budget since applying PGD to FastProtect tends to decrease invisibility. Hence, the baseline budget is adjusted to match the invisibility level of FastProtect + PGD. Surprisingly, our method serves as a superior initial checkpoint for the iterative optimization techniques. For example, the baseline, requires 100 steps to match our initial results, whereas FastProtect + PGD converges in just 25 steps with a significantly higher protection efficacy.
@@ 5 Conclusion
Acknowledgment. This work was partly supported by NAVER WEBTOON and by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT) under the Artificial Intelligence Convergence Innovation Human Resources Development, Inha University (No.RS-2022-00155915) and under the Leading Generative AI Human Resources Development (IITP-2025-RS-2024-00360227).
@@ Appendix A Implementation Details
Training. We set the number of perturbations $K$ in MoP to four. When calculating MLP loss in Eq. 5, we use intermediate features from [“down_1”, “down_2”, “down_3”, “mid_0”] layers in SD VAE, and set $\lambda$ to $3.5\times 10^{-5}$. We train FastProtect on a single A100 80GB GPU in an end-to-end manner except the assignment function $\mathcal{A}$. We use a batch size of 16, using the Adam optimizer [22] for 40k steps with learning rate of 0.0002, and betas of (0.5, 0.99). When training, the required VRAM and compute time are around 50GB and 12 hours, respectively. Note that since we update the perturbations with Adam instead of the direct optimization through PGD, we remove the minus sign in the training loss (Eq. 5) as: $\mathcal{L}_{T}(\mathbf{x})=||\mathbf{z}-\mathbf{z}_{y}||^{2}_{2}+\frac{\lambda}{L}\sum_{l=1}^{L}||\mathcal{F}^{l}-\mathcal{F}^{l}_{y}||^{2}_{2}.$ We utilize three patterned target images, each representing low, mid, and high pattern repetition, as illustrated in Figure 8. To implement adaptive targeted protection, for each protection strength (budget) $\eta$, we obtain three MoP models (low, mid, and high), each corresponding to one of the target images. The training algorithm is detailed in Algorithm 1; note that for simplicity, it is demonstrated with a batch size of one, but increasing the batch size is straightforward.
@@ Appendix B Related Work
Universal Adversarial Perturbation (UAP). Moosavi-Dezfooli et al. [36] demonstrated that a single perturbation could be used to attack various images, in contrast to the traditional image-specific perturbations. This approach is significantly more efficient in terms of computation time compared to per-instance adversarial attacks because it does not require iterative optimization for each individual image. Subsequently, Hayes and Danezis [14] presented a method for generating UAPs using generative models, showing improved efficiency and effectiveness over previous methods. Similarly, Mopuri et al. [38] utilized generative networks to create adversarial examples. In addition, there have been various attempts to find UAPs tailored to the image recognition tasks [5, 29, 37, 39]. However, these methods are not aligned with and cannot be adapted to the image protection task. Instead, in our study, we designed MoP with a focus on creating a protection framework that is both fast and maintains high performance. Our approach is specialized for image protection, differing from the general UAP-related methodologies introduced in adversarial attack field.
@@
AdvDM [28] 1210s /   35s 0.129 / 199.5 0.152 / 303.7 0.117 / 346.8 0.194 / 196.5
PhotoGuard [46]   370s /    7s 0.184 / 211.9 0.258 / 371.1 0.165 / 377.4 0.221 / 227.6
Anti-DB [51] 7278s / 225s 0.164 / 197.4 0.193 / 317.4 0.143 / 343.9 0.231 / 184.6
Mist [27] 1440s /   40s 0.185 / 217.2 0.259 / 365.8 0.166 / 386.2 0.223 / 223.7
Impasto [1]   830s /   19s 0.131 / 188.9 0.198 / 298.4 0.123 / 352.4 0.179 / 201.1
SDST [56] 1410s /   24s 0.170 / 198.0 0.244 / 302.1 0.152 / 354.1 0.199 / 196.7
@@
Invisible Adversarial Perturbation. Achieving high invisibility in adversarial attacks has garnered significant research interest. Several methods focus on restricting the regions of perturbations. Some approaches use the $L_{0}$ norm to generate sparse perturbations [35, 7], while others confine perturbations to small, salient areas to reduce visual distortions [8]. Various studies have targeted specific elements such as low-frequency components [13, 32] and have utilized advanced constraints like color components [59, 47] or quality assessments [53]. Additionally, Luo et al. [31] explored techniques for creating invisible and robust adversarial examples based on the human visual system. More recently, Laidlaw et al. [25] investigated perceptual adversarial robustness and adopted LPIPS [58] in adversarial perturbation optimization.
Although invisibility has been highly emphasized in adversarial attacks, it is less explored in the image protection task. Many methods enhance invisibility implicitly by minimizing the budget used, but this often leads to a trade-off with protection efficacy, resulting in decreased performance. Ahn et al. [1] constrain perturbations based on the human visual system using various modules during the optimization process. While effective, they require many ad-hoc components. In contrast, FastProtect uses LPIPS to create perceptual maps and apply masking, which is a simpler and more streamlined approach. This differs from Zhang et al. [58], who use LPIPS in the optimization process.
@@ Appendix C Experimental Setups
Datasets. We utilize the Object, Face, Painting, and Cartoon domains in both the training and benchmark datasets. When constructing the training dataset, we randomly sample 20k images from ImageNet [9] for the Object domain, resizing all images to $512^{2}$. No augmentation is applied. For the Face domain, we use 20k images randomly sampled from the FFHQ dataset [19]. The images, originally $1024^{2}$ and face-aligned [20], are all downscaled to $512^{2}$, without further augmentations. The Painting dataset is created by randomly sampling 20k images from the WikiArt dataset [45], resizing them to $512^{2}$. We filter out images with resolutions outside the range of 512 to 2048-px. For the Cartoon dataset, we use Webtoon artworks published on NAVER Webtoon, sampling 20k images only the works with permission from the original creators for research purposes. Similar to the Painting dataset, we apply resolution filtering and resize the images to $512^{2}$.
To make the benchmark dataset, we select 20 objects from the personalization subject dataset proposed by Ruiz et al. [44] for the Object domain. Each object consists of 5-6 images. When we conduct the main comparison (Table 2), we use images resized to $512^{2}$, and for the arbitrary image analysis (Table 4), we apply protection to the original resolution images. For the Face domain, we followe Van Le et al. [51] and randomly sample 20 identities from VGGFace2 [4], each identity consisting of 12 images resized to $512^{2}$. The Painting dataset is compiled by randomly sampling (but not overlap to the training dataset) 20 artists from the WikiArt dataset, each artist represented by 10 artworks, all resized to $512^{2}$. For the Cartoon domain, we randomly sample 20 Webtoon works (but also no overlap to the training dataset) from the Webtoon dataset. In this dataset, we only use facial images of major cartoon characters by cropping and aligning all the images. This reflects the common practice in cartoons and illustrations where characters are the main focus to both readers and artists. In the main comparison we use $512^{2}$ images, while the arbitrary analysis maintains the original resolution.
@@
AdvDM [28] 0.197 0.362 0.540 220.0 0.458 2.139
PhotoGuard [46] 0.203 0.347 0.575 223.0 0.506 2.396
Mist [27] 0.185 0.322 0.578 217.2 0.523 2.328
SDST [56] 0.242 0.402 0.575 219.2 0.542 2.442
Anti-DB [51] 0.239 0.413 0.510 214.4 0.439 2.148
Impasto [1] 0.201 0.378 0.588 213.8 0.473 2.183
@@
AdvDM [28] 0.173 0.338 0.528 303.8 0.466 2.493
PhotoGuard [46] 0.189 0.357 0.557 308.7 0.630 3.113
Mist [27] 0.154 0.310 0.560 307.5 0.667 3.131
SDST [56] 0.244 0.431 0.561 302.1 0.629 3.019
Anti-DB [51] 0.162 0.309 0.543 301.4 0.484 2.411
Impasto [1] 0.198 0.388 0.562 298.4 0.565 2.939
@@
Baselines. FastProtect is compared with existing diffusion-based mimicry protection frameworks: AdvDM [28], PhotoGuard [46], Anti-DreamBooth (Anti-DB) [51], Mist [27], Impasto [1], and SDST [56]. We follow the official settings for all the baselines by using their official codes. However, for PhotoGuard, we use the target image of Mist [27] by following the protocols of Xue et al. [56] while we use the Impasto’s target image for the Impasto.
Evaluation. For the primary evaluation assessment, we use DISTS [10] to measure the invisibility of the protected image and FID [15] to evaluate the protection efficacy of the personalized diffusion methods. In addition, in here, we also compare the protection frameworks using other metrics: To evaluate invisibility, we include LPIPSVGG [58] and AHIQ [26]. Note that LPIPSVGG measures the perceptual distance in the VGG feature space, which is different from the feature space used in our perceptual map creation module (i.e. AlexNet). For evaluating protection efficacy, we use TOPIQ-NR [6] and QAlign [54].
When preparing mimicry outputs, we personalize the diffusion models using the protected image with LoRA [17] by default, employing default settings for all the personalization methods. To generate outputs, we use different inference prompts than those used during training to simulate black-box caption scenarios [51]. For example, we train the diffusion models with LoRA using captions of “A painting in <sks>  style” prompt and perform inference with “A painting of a house in <sks>  style”.